OpenStudy (anonymous):

Can someone PLEASE help me, Step by step... Sara and Elizabeth went shopping together at their favorite store during a sale. They both bought some dress shirts that cost $25.00 a piece t-shirts that cost $10.50 a piece. If sara bought 3 more dress shirts than elizabeth and they both bought 4 t-shirts, how many dress shirts did sarah buy if together they spent a total of $309.00 (before taxes) on these items?

OpenStudy (anonymous):

first I would say 309.00-75.00(the price of the extra dress shirts)= 234.00/2=117.00 and then add liz's 75.00 back in 117.00+75.00=192.00 Then to check your answer add 192.00+117.00=309.00

OpenStudy (anonymous):

was that any help???

OpenStudy (anonymous):

Yes, thank you. But i have to set up and equation with variables.

OpenStudy (anonymous):

oooohhhhhh like you make liz X and sara x2?

OpenStudy (anonymous):

No, where you make sara x and liz y